# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v2

@@ Appendix G Fixation probability in time-switching environments under weak bottlenecks
Interestingly however, Fig. 4(c) illustrates the almost unnoticeable dependence of $\Phi^{\rm G}(\nu,m)$ on the specific spatial structure. This is due to the overall similar impact of the factor $E_{\rm G}(i)/q_{G}$ for the various graphs. While differences arise when $m$ varies at fixed $\nu$ due to large variations in the timescales of the competition dynamics, varying spatial structure produces small changes in these timescales, and as such leads to only unnoticeable changes in $\Phi$.
@@ Appendix H Asymmetric dichotomous Markov noise & environmental bias
which is the same expression as Eq. (6), but now driven by the aDMN $\xi(t)$ [16, 17, 18, 19, 21, 22, 32]. The carrying capacity switches back and forth from $K_{+}$ (mild environment, $\xi=1$) to $K_{-}<K_{+}$ (harsh environment, $\xi=-1$) at rates $\nu_{\pm}=\nu(1\mp\delta)$ according to
@@
The regime of strong bottlenecks is dominated by the interplay between $M/W$ competition in the mild state ($K=K_{+}$) and deme extinction in the harsh environmental state ($K=K_{-}$), occurring in time $\theta_{E}\equiv\theta_{E}(K_{-},\Omega)$. In this regime, the near-optimal conditions for the removal of the mutant strain can be obtained as under symmetric switching (given by Eq. (27)) and now read:
@@ Appendix I Simulation methods & plots
While not exact like other simulation methods (e.g. the Gillespie algorithm [97]) due to time-discretisation, the MC algorithm used here improves on computational efficiency making simulations with larger numbers of cells on the metapopulation that run for long times feasible and straightforward to implement. The questions of computational efficiency and tractability are particular critical in the context of this work in which we study fixation and extinction of spatially arranged populations, a notoriously computationally demanding problem. In this context MC algorithms similar to the one used are useful tools to investigate the properties of spatially extended systems; see, e.g., Ref. [98].
In the case of symmetric environmental switching, the MC algorithm that we have employed can be described as follows: The graph, consisting of the spatially arranged $\Omega$ deme forming the metapopulation, is initialised by randomly picking an initial value for the carrying capacity $K(0)=\{K_{-},K_{+}\}$ with equal probability (in the constant environment case $K(0)=K$ with probability 1), populating a single deme with $K(0)$ $M$-cells, and the remaining $\Omega-1$ of the demes with $K(0)$ $W$-cells. Time is then discretised in units of Monte Carlo steps (MCS) whereby in each MCS we perform $2N$ birth and death events, where $N$ is measured at the start of the MCS. We choose $2N$ as we typically have $N$ births and $N$ deaths per unit time according to the transition rates Eq. (2) and Eq. (4) when summing over all cells on the metapopulation. Therefore, our units of time are consistent between the theoretical model and the Monte Carlo simulation. Each of the $2N$ events in an MCS occur sequentially and the rates are updated for the next event in the MCS. The type of each of the $2N$ events within an MCS is selected based on the rates of the events, where a higher rate means a larger probability of that event being selected. Concretely, the following steps for a single MCS occur:
  • Check if an environmental switch, occurring with rate $\nu$, occurs on the metapopulation. The rate of reaction of birth/death/migration events on the entire metapopulation is $N(1+m+N/K(t))$. Therefore, the probability of an environmental switch is given by $\nu/(N(1+m+N/K(t)+\nu)$.
@@
  • The above steps are repeated until $2N$ births and deaths events are performed.
In all simulations, each realisation is simulated until fixation, tracking when fixation occurs and which species fixates. In simulations for Figs. 2(e,f), 6, 10, and 13 the simulation then continues until metapopulation extinction or a large fixed time $T$ (here we set $T=10^{5}$). For a given set of parameters, if extinction does not occur in any realisation, the extinction time is not recorded. The data is averaged to obtain the fixation probability, the mean fixation time, and the mean time to extinction where applicable. Furthermore, we record the standard error on the mean for each quantity. In Figs. 2, 4, and 12, $10^{3}$ realisations are ran for each set of parameters, and these data are plotted with the standard error on the mean shows as error bars. In the heatmaps, each data point corresponds to the average value for $10^{3}$ simulations at that point in parameter space. The standard error on the mean is not plotted for the heatmaps.

### v3

@@ Appendix G Fixation probability in time-switching environments under weak bottlenecks
Interestingly however, Fig. 4(c) illustrates the almost unnoticeable dependence of $\Phi^{\rm G}(\nu,m)$ on the specific spatial structure. This is due to the overall similar impact of the factor $E_{\rm G}(i)/q_{G}$ for the various graphs. While differences arise when $m$ varies at fixed $\nu$ due to large variations in the timescales of the competition dynamics, varying spatial structure produces small changes in these timescales, and as such leads to only unnoticeable changes in $\Phi^{\rm G}$.
@@ Appendix H Asymmetric dichotomous Markov noise & environmental bias
which is the same expression as Eq. (6), but now driven by the aDMN $\xi(t)$ [16, 17, 18, 19, 21, 22, 32]. The carrying capacity switches back and forth between $K_{+}$ (mild environment, $\xi=+1$) and $K_{-}<K_{+}$ (harsh environment, $\xi=-1$) at rates $\nu_{\pm}=\nu(1\mp\delta)$ according to
@@
The regime of strong bottlenecks is dominated by the interplay between $M/W$ competition in the mild state ($K=K_{+}$) and deme extinction in the harsh environmental state ($K=K_{-}$), occurring in time $\theta_{E}\equiv\theta_{E}(K_{-},\Omega)$. In this regime, the near-optimal conditions for the removal of the mutant strain can be obtained as under symmetric switching (given by Eq. (27)) and read
@@ Appendix I Simulation methods & plots
While not exact-like other simulation methods (e.g. the Gillespie algorithm [97]) due to time-discretisation, the MC algorithm used here improves on computational efficiency making simulations with larger numbers of cells on the metapopulation that run for long times feasible and straightforward to implement. The questions of computational efficiency and tractability are particularly critical in the context of this work in which we study fixation and extinction of spatially arranged populations, a notoriously computationally demanding problem. In this context MC algorithms similar to the one used are useful tools to investigate the properties of spatially extended systems; see, for example, Ref. [98].
In the case of symmetric environmental switching, the MC algorithm that we have employed can be described as follows: The graph, consisting of the spatially arranged $\Omega$ deme forming the metapopulation, is initialised by randomly picking an initial value for the carrying capacity $K(0)=\{K_{-},K_{+}\}$ with equal probability (in the constant environment case $K(0)=K$ with probability 1), populating a single deme with $K(0)$ $M$-cells, and the remaining $\Omega-1$ of the demes with $K(0)$ $W$-cells. Time is then discretised in units of Monte Carlo steps (MCS) whereby in each MCS we perform $2N$ birth/death events, where $N$ is measured at the start of the MCS. We choose $2N$ as we typically have $N$ births and $N$ deaths per unit time according to the transition rates Eq. (2) and Eq. (4) when summing over all cells on the metapopulation. Therefore, our units of time are consistent between the theoretical model and the Monte Carlo simulation. We emphasise that migration events and environmental switches do not contribute to the $2N$ events comprising an MCS. Therefore, $2N$ birth/death events are performed in a single MCS, and the expected number of migrations and environmental switches in the MCS corresponds to the migration and switching events expected in one unit of time. Each of the $2N$ events in an MCS occurs sequentially and the rates are updated for the next event in the MCS. The type of each event in an MCS is selected sequentially based on the rates of the events, where a higher rate means a larger probability of that event being selected. Concretely, the following steps for a single MCS occur:
  • Check if an environmental switch, occurring with rate $\nu$, occurs on the metapopulation. The rate of reaction of birth/death/migration events on the entire metapopulation is $N(1+m+N/K(t))$. Therefore, the probability of an environmental switch is given by $\nu/(N(1+m+N/K(t))+\nu)$.
@@
  • The above steps are repeated until $2N$ birth/death events are performed.
Since each event is selected based on the current propensities of the system, and those are subsequently updated before the selection of the next event. The probability of events in our MC algorithm is thus statistically equivalent to that of a simulation generating the statistically correct sample paths (e.g. Gillespie [97]). The approximation of this algorithm is to assign each MCS to $2N$ birth/death events with $N$ defined at the start of the MCS, unlike a statistically exact simulation algorithm where each event time is drawn individually.
In all simulations, each realisation is simulated until fixation, tracking when fixation occurs and which species fixates. In simulations for Figs. 2(b), 6, 10, and 13 the simulation then continues until metapopulation extinction or a large fixed time $T$ (here we set $T=10^{5}$). For a given set of parameters, if extinction does not occur in any realisation, the extinction time is not recorded. The data is averaged to obtain the fixation probability, the mean fixation time, and the mean time to extinction where applicable. Furthermore, we record the standard error on the mean for each quantity. In Figs. 2, 4, and 12, $10^{3}$ realisations are ran for each set of parameters, and these data are plotted with the standard error on the mean shows as error bars. In the heatmaps, each data point corresponds to the average value for $10^{3}$ simulations at that point in parameter space. The standard error on the mean is not plotted for the heatmaps.